# A Multimodal Manufacturing Safety Chatbot: Knowledge Base Design, Benchmark Development, and Evaluation of Multiple RAG Approaches

### v1

@@ 6.3 Phase 6: Human Evaluation
While automated evaluations provide valuable initial insights, they can only go so far in identifying the best-performing retrieval–generation pipeline. For example, metrics such as LLM-as-judge may be imperfect for distinguishing accuracy across competing solutions. To complement these quantitative results and select a single final pipeline for potential deployment, we conducted a human expert evaluation.
In the first stage, we evaluated all 24 pipelines on the full UR5 Cobot benchmark. Performance was assessed primarily on accuracy. Latency was considered as a secondary criterion, and cost was treated as an acceptability threshold rather than a ranking factor. When pipelines had no practical differences in accuracy or latency, we relied on their accuracy ranking to guide the selection of candidates. Based on this full-benchmark evaluation, we narrowed the original set of 24 pipelines down to 2.
After this narrowing step, we drew a random sample of ten questions from the UR5 Cobot benchmark for human evaluation. For each sampled question, we retrieved the corresponding answers generated by the two selected pipelines. These ten questions and answers were then presented to expert evaluators through a purpose-built web interface that enabled side-by-side comparison of model outputs. Figure 4 shows this evaluation interface. This setup allowed us to distribute a structured survey to domain specialists in manufacturing, occupational safety, and human factors.
Figure 4: A screenshot of the interface used for human evaluation.
@@ 6.4.1 Accuracy
Finally, we investigated the main effect of retrieval approach on answer accuracy. Cochran’s $Q$ test indicated a significant effect of retrieval method, Q(5)=103.24,p<0.001Q(5)=103.24,\ p<0.001. Averaged across models and retrieval depths, the highest proportions of correct answers were observed for the OpenAI Keyword and OpenAI Semantic approaches (both $0.77$), followed by BM25 ($0.71$), Vanilla ($0.68$), Graph MMR ($0.62$), and Graph Eager ($0.50$). Post-hoc pairwise McNemar tests with Holm adjustment, focusing on the three best-performing methods, revealed no significant differences among OpenAI Keyword, OpenAI Semantic, and BM25 (all adjusted $p=0.2$), indicating statistically equivalent performance at the top end. Both OpenAI-based methods, however, significantly outperformed all remaining baselines, namely Graph Eager, Graph MMR, and Vanilla (all adjusted $p<0.05$), highlighting the advantage of these learned retrieval strategies for maximizing response correctness.
It is important to note that these results characterize average (main) effects of each factor. That is, we ask questions such as “Is mini more accurate than nano on average?” or “Which retrieval approach performs best on average?” rather than examining the full pattern of interactions (for example, whether certain retrieval methods benefit mini more than nano, or whether top k = $7$ helps some methods but not others). Analyses at the level of specific model–method–depth combinations, such as those visualized in Figure 5A, would be required to fully characterize such interaction effects.
@@ 6.4.3 Cost
To complement the accuracy and latency analyses, we evaluated how each of the model size, retrieval depth, and retrieval method affected the monetary cost per request (Figure 5C). Costs were computed from input and output token usage under the October 2025 pricing for gpt-5-mini-2025-08-07 and gpt-5-nano-2025-08-07 [openai2025pricing].
@@ 7 Concluding Remarks
This study introduced a multimodal, retrieval-augmented chatbot framework designed to advance safety training and hazard recognition within Industry 5.0 manufacturing environments. Grounded in a rigorous Design Science Research (DSR) methodology, the work systematically addressed three critical requirements for next-generation safety systems, namely, accuracy, latency, and cost, through an experimentally validated approach. The final system integrates text, image, and voice modalities with multiple retrieval strategies to deliver contextually grounded, real-time safety guidance. Beyond the artifact itself, a secondary contribution of our work is a publicly available benchmark that can help to evaluate future AI solutions in manufacturing settings.
We argue that our empirical results are of great value to practitioners, as they demonstrate how different retrieval strategies and model configurations can exert a significant influence on system performance. In particular, our findings highlight the multi-objective nature of RAG configuration optimization, where trade-offs among correctness, speed, and expense must be balanced. This leads to interesting future research directions, where one can explore Pareto-based optimization frameworks to systematically identify non-dominated solutions that best balance accuracy, latency, and cost given a manufacturing facility’s needs and budget. Our results with a local RAG approach suggest another promising avenue for future work. In particular, by fine-tuning lightweight local LLMs and combining them with efficient, on-premise RAG pipelines, one can potentially reduce both operational costs and response latency while preserving accuracy. These directions can further enhance our contribution of a framework and empirical foundation for next-generation AI-driven safety training systems, which marks a step toward human–AI collaboration that enhances safety, efficiency, and resilience in smart manufacturing environments.
@@
All code and evaluation benchmarks used in this study are publicly available at https://github.com/fmegahed/safety_rag_evaluation.
@@ Acknowledgments
This work was funded by the Ohio Bureau of Workers’ Compensation Worker Safety Innovation Center (WSIC) Grant #WSIC26-250206-009 and by the Raymond E. Glos Professorship and the Dinesh & Ila Paliwal Innovation Chair at Miami University. The authors also acknowledge the Farmer School of Business for providing the computing infrastructure used to host the chatbot developed in this research. The authors are also thankful to Engineered Profiles, Yamaha Motor Company of North America, and MeetKai, Inc. for their feedback on our benchmark Q&A datasets.

### v2

@@ 6.3 Phase 6: Human Evaluation
While automated evaluations provide valuable initial insights, they can only go so far in identifying the best retrieval–generation pipeline. For example, metrics such as LLM-as-judge may be imperfect for differentiating accuracy across solutions. To complement these results and select a single final pipeline, we conducted a human expert evaluation.
In the first stage, we evaluated all 24 pipelines on the full UR5 Cobot benchmark. Performance was assessed primarily on accuracy. Latency was considered as a secondary criterion, and cost was treated as an acceptability threshold rather than a ranking factor. When pipelines had no practical differences in accuracy or latency, we relied on their accuracy ranking to guide the selection of candidates. Based on this evaluation plan, we narrowed the original set of 24 pipelines down to 2.
After this narrowing step, we drew a random sample of ten questions from the UR5 Cobot benchmark for human evaluation. For each sampled question, we retrieved the corresponding answers generated by the two selected pipelines. These ten questions and answers were then presented to expert evaluators through a purpose-built web interface that enabled blind side-by-side comparison of model outputs. Figure 4 shows this evaluation interface. This setup allowed us to distribute a structured survey to domain specialists in manufacturing, occupational safety, and human factors. All evaluators were safety engineers and academic researchers; we did not include front-line operators in this study. A planned future study includes assessing how performance and reliance change for less technically literate user groups, including new operators.
Figure 4: A screenshot of the interface used for human evaluation for blind side-by-side comparison of model outputs.
@@ 6.4.1 Accuracy
Finally, we investigated the main effect of retrieval approach on answer accuracy. Cochran’s $Q$ test indicated a significant effect of retrieval method, Q(5)=103.24,p<0.001Q(5)=103.24,\ p<0.001. Averaged across models and retrieval depths, the highest proportions of correct answers were observed for the OpenAI Keyword and OpenAI Semantic approaches (both $0.77$), followed by BM25 ($0.71$), Vanilla ($0.68$), Graph MMR ($0.62$), and Graph Eager ($0.50$). Post-hoc pairwise McNemar tests with Holm adjustment, focusing on the three best-performing methods, revealed no significant differences among OpenAI Keyword, OpenAI Semantic, and BM25 (all adjusted $p=0.2$), indicating statistically equivalent performance at the top end. Both OpenAI-based methods, however, significantly outperformed all remaining baselines, i.e., Graph Eager, Graph MMR, and Vanilla (all adjusted $p<0.05$), highlighting the advantage of OpenAI’s retrieval strategies for maximizing response correctness.
These results characterize average (main) effects of each factor. That is, we ask questions such as “Is mini more accurate than nano on average?” or “Which retrieval approach performs best on average?” rather than examining the full pattern of interactions (for example, whether certain retrieval methods benefit mini more than nano, or whether top k = $7$ helps some methods but not others). Analyses at the level of specific model–method–depth combinations, e.g., those in Figure 5A, would be required to fully characterize such interaction effects.
@@ 6.4.2 Latency
It is important to note that our latency measurements reflect end-to-end time to generate the entire response, rather than time-to-first-token. As a result, these values can overestimate perceived latency in practice, since users typically begin reading or listening while the response is still being generated. More importantly, the chatbot is intended for informational safety guidance (e.g., training and decision support) and is not designed for time-critical protective actions such as emergency stop activation or lock-out/tag-out execution, for which latencies on the order of seconds (or less) are required.
@@ 6.4.3 Cost
To complement the accuracy and latency analyses, we evaluated how each of the model size, retrieval depth, and retrieval method affected the monetary cost per request (Figure 5C). Costs were computed from input and output token usage under the October 2025 pricing for gpt-5-mini-2025-08-07 and gpt-5-nano-2025-08-07.
@@ 7 Concluding Remarks
This study introduced a multimodal, RAG-based chatbot framework designed to advance safety training and hazard recognition within Industry 5.0 manufacturing environments. Grounded in a rigorous Design Science Research methodology, the work systematically addressed three critical requirements for next-generation safety systems, namely, accuracy, latency, and cost, through an experimentally validated approach. The final system integrates text, image, and voice modalities with multiple retrieval strategies to deliver contextually grounded, real-time safety guidance. Beyond the artifact, a secondary contribution of our work is a publicly available benchmark that can help to evaluate future AI manufacturing solutions.
We argue that our empirical results are of great value to practitioners, as they demonstrate how different retrieval strategies and model configurations can exert a significant influence on system performance. In particular, our findings highlight the multi-objective nature of RAG configuration optimization, where trade-offs among correctness, speed, and expense must be balanced. This leads to interesting future research directions, where one can explore Pareto-based optimization frameworks to systematically identify non-dominated solutions that best balance accuracy, latency, and cost given a manufacturing facility’s needs and budget. Our results with a local RAG approach suggest another promising avenue for future work. In particular, by fine-tuning lightweight local LLMs and combining them with efficient, on-premise RAG pipelines, one can potentially reduce both operational costs and response latency while preserving accuracy. Finally, given the safety-critical context, future work should evaluate trust calibration by measuring over-reliance (accepting incorrect advice) and under-reliance (ignoring correct guidance) in realistic user studies. Such studies can also test behavioral interventions to better align reliance with system reliability. These directions can further enhance our contribution of a framework and empirical foundation for next-generation AI-driven safety training systems, which marks a step toward human–AI collaboration that enhances safety, efficiency, and resilience in smart manufacturing environments.
@@
Code and evaluation benchmarks are available at https://github.com/fmegahed/safety_rag_evaluation.
@@ Acknowledgments
This work was funded by the Ohio Bureau of Workers’ Compensation Worker Safety Innovation Center (WSIC) Grant #WSIC26-250206-009. Dr. Megahed and Carvalho also received support from the Raymond E. Glos Professorship and the Dinesh & Ila Paliwal Innovation Chair. The authors acknowledge the Farmer School of Business for hosting our chatbot. The authors also appreciate the valuable feedback provided by Engineered Profiles, Yamaha Motor Company of North America, and MeetKai, Inc. on our benchmark Q&A datasets.